# Adaptive Layer Splitting for Wireless LLM Inference in Edge Computing: A Model-Based Reinforcement Learning Approach

### v4

@@ II-A Edge-Enhanced LLM Deployment
In the realm of LLMs, [43, 44] provide a comprehensive overview of the challenges faced by LLMs in cloud-based settings, particularly emphasizing the constraints related to data privacy and processing efficiency. [45] propose the concept of edge computing as a viable solution. Subsequently, [21] propose an LLM cloud-edge collaboration framework, which utilizes small-scale models deployed at edge nodes to enhance the generative capabilities of cloud-based LLMs. Additionally, [22] emphasize the importance of optimizing LLM deployment at 6G edges but do not provide specific methodologies for leveraging edge computing capabilities. [46] propose to use LLMs as offline compilers to meet low-latency requirements in edge computing. However, these approaches do not fully address the computational constraints of edge nodes. Our research bridges this gap by introducing a dynamical layer splitting framework to alleviate these limitations.
@@ III-A System Model
where $\theta_{\text{UE}}$ denotes the parameters of layers in $L_{U}(\cdot)$. This procedure inevitably incurs a certain computational load on UEs, typically measured in floating-point operations per second (FLOPs). Typically, for the layer $1$ with an input sequence of length $d_{\text{in}}$ with hidden dimension $d_{\text{mid}}$, utilizing a multi-head self-attention mechanism with $\kappa$ heads, the computations for the involved multi-head attention mechanism and feed-forward network components can be obtained as $\text{FLOPs}(L_{1})=\frac{3d_{\text{in}}d_{\text{mid}}^{2}}{\kappa}+\frac{2d_{\text{in}}^{2}d_{\text{mid}}}{\kappa}+9d_{\text{in}}d_{\text{mid}}^{2}$. Thus, the computational load on the UE can be approximated as
$C_{\text{UE}}(p)$ $=\sum_{i=1}^{p}\text{FLOPs}(L_{i})$ (2)
$=p\cdot\big{(}\frac{3d_{\text{in}}d_{\text{mid}}^{2}}{\kappa}+\frac{2d_{\text{in}}^{2}d_{\text{mid}}}{\kappa}+9d_{\text{in}}d_{\text{mid}}^{2}\big{)}.$
@@ III-B Problem Formulation
On the other hand, high computational load on UEs would lead to increased latency and energy consumption, further undermining the benefits of deploying LLMs in edge environments. As indicated in (III-A), the computational load on the UE is approximately proportional to the number of layers processed locally. This proportionality yields a contradicting phenomenon that an earlier model splitting worsens the inference performance but ameliorates the computational cost at the computation-limited UE. In other words, in order to minimize the overall system PPL while effectively balancing the computational load on the UE, the problem turns to identifying the optimal splitting point $p$ under volatile channel conditions, that is,
@@
where $\text{PPL}(p,\sigma,m)$ quantifies the LLM’p inference performance, taking into account the splitting point $p$, the noise intensity $\sigma$, and the Nakagami-$m$ fading shape parameter $m$, which directly affects the packet loss probability. Besides, the weight $\lambda$ balances the trade-off between the inference performance and computational load.
@@ IV-A The Markov Decision Process
The splitting point adjustment under volatile channels can be formalized as a Markov Decision Process (MDP), consisting of a tuple $\langle\mathcal{S},\mathcal{A},P,R\rangle$. In particular, as illustrated in Fig. 4, the state space $\mathcal{S}$ encompasses the key factors such as the noise intensity $\sigma$, the Nakagami-$m$ fading shape $m$ and the current splitting point $p$, namely, $s=\{\sigma,m,p\}\in\mathcal{S}$. The action space $\mathcal{A}$ is designed to accommodate a range of possible adjustments to the splitting point, allowing for both fine-grained and more substantial modifications. Specifically, $\mathcal{A}$ includes actions such as moving the splitting point upward or downward by $u$ layers, where $n$ can take values such as $1$, $2$, $3$, and so forth, or maintaining the current position. Mathematically, this can be expressed as $a\in\mathcal{A}=\{-u,\dots,-1,0,+1,\dots,+u\}$. This generalized action space provides the RL agent with the flexibility to optimize the splitting strategy dynamically in response to varying channel environments. For a time-step $t$, given an action $a_{t}$ under state $s_{t}$, the environment state will transit to the next state $s_{t+1}$ following the transition probability $P$, which is contingent on the selected action and real-time channel conditions. Meanwhile, a reward can be obtained as
@@ IV-C The Reward Surrogate Model for Faster RL
$\tilde{R}(s_{t},a_{t})=-\left(\widetilde{\text{PPL}}(p_{t},\sigma_{t},m_{t},\theta_{\text{surr}})+\lambda\cdot C_{\text{UE}}(p_{t})\right),$ (15)
where $\widetilde{\text{PPL}}(p_{t},\sigma_{t},m_{t},\theta_{\text{surr}})$ represents the estimated PPL provided by the surrogate model. This reformulation significantly reduces the training burden by replacing the direct LLM inference with an efficient approximation, enabling more rapid policy evaluation and iteration of an RL policy. From a theoretical standpoint [64], the surrogate model effectively reduces the variance in reward estimation by providing a smoothed approximation of the true reward landscape. This smoothing is particularly advantageous in high-dimensional action spaces, where small perturbations in actions could lead to large fluctuations in PPL if calculated directly.
@@ V-A Experimental Setup
To reduce the computational cost of evaluating the LLM’s performance at each step, by collecting $9,718$ pieces of practical records, we derive a reward surrogate model as in Section IV-C. Our result shows that an MLP yields a test loss of $0.00548$ in MSE and $0.050$ in mean absolute error (MAE), thus providing sufficient accuracy. Therefore, we use this MLP-based reward surrogate model to accelerate the evaluation process.

### v5

@@ II-A Edge-Enhanced LLM Deployment
In the realm of LLMs, [43, 44] provide a comprehensive overview of the challenges faced by LLMs in cloud-based settings, particularly emphasizing the constraints related to data privacy and processing efficiency. [45] proposes the concept of edge computing as a viable solution. Subsequently, [21] proposes an LLM cloud-edge collaboration framework, which utilizes small-scale models deployed at edge nodes to enhance the generative capabilities of cloud-based LLMs. Additionally, [22] emphasizes the importance of optimizing LLM deployment at 6G edges but does not provide specific methodologies for leveraging edge computing capabilities. [46] proposes to use LLMs as offline compilers to meet low-latency requirements in edge computing. However, these approaches do not fully address the computational constraints of edge nodes. Our research bridges this gap by introducing a dynamical layer splitting framework to alleviate these limitations.
@@ III-A System Model
where $\theta_{\text{UE}}$ denotes the parameters of layers in $L_{U}(\cdot)$. This procedure inevitably incurs a certain computational load on UEs, typically measured in floating-point operations per second (FLOPs). Typically, for the layer $1$ with an input sequence of length $d_{\text{in}}$ and hidden dimension $d_{\text{mid}}$, utilizing a multi-head self-attention mechanism with $\kappa$ heads, the computations for the involved multi-head attention mechanism and feed-forward network components can be obtained as $\text{FLOPs}(L_{1})=\frac{3d_{\text{in}}d_{\text{mid}}^{2}}{\kappa}+\frac{2d_{\text{in}}^{2}d_{\text{mid}}}{\kappa}+9d_{\text{in}}d_{\text{mid}}^{2}$. Thus, the computational load on the UE can be approximated as
$C_{\text{UE}}(p)$ $=\sum_{i=1}^{p}\text{FLOPs}(L_{i})$
$=p\cdot\big{(}\frac{3d_{\text{in}}d_{\text{mid}}^{2}}{\kappa}+\frac{2d_{\text{in}}^{2}d_{\text{mid}}}{\kappa}+9d_{\text{in}}d_{\text{mid}}^{2}\big{)}.$ (2)
@@ III-B Problem Formulation
On the other hand, high computational load on UEs would lead to increased latency and energy consumption, further undermining the benefits of deploying LLMs in edge environments. As indicated in (2), the computational load on the UE is approximately proportional to the number of layers processed locally. This proportionality yields a contradicting phenomenon that an earlier model splitting worsens the inference performance but ameliorates the computational cost at the computation-limited UE. In other words, in order to minimize the overall system PPL while effectively balancing the computational load on the UE, the problem turns to identifying the optimal splitting point $p$ under volatile channel conditions, that is,
@@
where $\text{PPL}(p;\sigma,m)$ quantifies the LLM’p inference performance, taking into account the splitting point $p$, the noise intensity $\sigma$, and the Nakagami-$m$ fading shape parameter $m$, which directly affects the packet loss probability. Besides, the weight $\lambda$ balances the trade-off between the inference performance and computational load.
@@ IV-A The Markov Decision Process
The splitting point adjustment under volatile channels can be formalized as a Markov Decision Process (MDP), consisting of a tuple $\langle\mathcal{S},\mathcal{A},P,R\rangle$. In particular, as illustrated in Fig. 4, the state space $\mathcal{S}$ encompasses the key factors such as the noise intensity $\sigma$, the Nakagami-$m$ fading shape $m$ and the current splitting point $p$, namely, $s=\{\sigma,m,p\}\in\mathcal{S}$. The action space $\mathcal{A}$ is designed to accommodate a range of possible adjustments to the splitting point, allowing for both fine-grained and more substantial modifications. Specifically, $\mathcal{A}$ includes actions such as moving the splitting point upward or downward by $u$ layers, where $u$ can take values such as $1$, $2$, $3$, and so forth, or maintaining the current position. Mathematically, this can be expressed as $a\in\mathcal{A}=\{-u,\dots,-1,0,+1,\dots,+u\}$. This generalized action space provides the RL agent with the flexibility to optimize the splitting strategy dynamically in response to varying channel environments. For a time-step $t$, given an action $a_{t}$ under state $s_{t}$, the environment state will transit to the next state $s_{t+1}$ following the transition probability $P$, which is contingent on the selected action and real-time channel conditions. Meanwhile, a reward can be obtained as
@@ IV-C The Reward Surrogate Model for Faster RL
$\tilde{R}(s_{t},a_{t})=-\left(\widetilde{\text{PPL}}(p_{t};\sigma_{t},m_{t},\theta_{\text{surr}})+\lambda\cdot C_{\text{UE}}(p_{t})\right),$ (15)
where $\widetilde{\text{PPL}}(p_{t};\sigma_{t},m_{t},\theta_{\text{surr}})$ represents the estimated PPL provided by the surrogate model. This reformulation significantly reduces the training burden by replacing the direct LLM inference with an efficient approximation, enabling more rapid policy evaluation and iteration of an RL policy. From a theoretical standpoint [64], the surrogate model effectively reduces the variance in reward estimation by providing a smoothed approximation of the true reward landscape. This smoothing is particularly advantageous in high-dimensional action spaces, where small perturbations in actions could lead to large fluctuations in PPL if calculated directly.
@@ V-A Experimental Setup
To reduce the computational cost of evaluating the LLM’s performance at each step, by collecting $9,718$ pieces of practical records, we derive a reward surrogate model as in Section IV-C. Our result shows that an MLP (Multi-Layer Perceptron) yields a test loss of $0.00548$ in MSE and $0.050$ in mean absolute error (MAE), thus providing sufficient accuracy. Therefore, we use this MLP-based reward surrogate model to accelerate the evaluation process.